Q ETV2 Libro 7.5 m Altura final 0.8 m Save A

You drop a 2.5 kg book, from rest, from a height of 7.5 m. A student receives it at a height of 0.8 m from the floor. If the zero potential energy reference level is the floor, the kinetic energy in the problem book just an instant before falling into the hand of the student who receives it below is:
Answer #1

Potential energy converted inti kinetic energy.

Potential energy = mg (hi - hf)

Where hi = Initial height = 7.5 m

hf = final height = 0.8 m

Potential energy = 2.5 × 9.81 × (7.5-0.8) = 164.3175 J

So, kinetic energy before he recieves the book = 164.3175 J